### v1

@@ 4.5 Datasets
For procedural error detection, various egocentric datasets have been introduced, covering scenarios such as cooking [damen2020epic][peddi2024captaincook4d], assembly [sener2022assembly101][liu2024ikea], outdoor tent pitching [jang2019epic], and specialized domains like chemical experiments [haneji2024egooops]. Several additional datasets are also prevalent in procedural learning research, as shown in Table 4.
Breakfast [kuehne2014language] includes 77 hours of video for 10 cooking activities by 52 participants in 18 kitchens. Each video averages 5.2 unique actions and 6.9 action segments, with 10% of actions repeated.
Inria [alayrac2016unsupervised] is an instructional dataset of 150 videos (30 per task) spanning five tasks, with 800,000 frames and comprehensive annotations.
YouCook2 [zhou2018towards] contains 2,000 untrimmed videos featuring 89 global cooking recipes, averaging 22 videos per recipe and encompassing diverse cooking styles.
COIN [tang2019coin] contains 11,827 YouTube videos across 180 tasks in 12 daily domains, totaling 476 hours and 46,354 annotated segments, with an average video length of 2.36 minutes.
HowTo100M [miech2019howto100m] includes 1.22 million instructional videos, covering 23,000+ visual tasks and 136 million clips, each with narrations and auto-downloaded subtitles from YouTube.
CrossTask [zhukov2019cross] consists of instructional videos for 83 tasks (e.g., making a bread, preparing a latte, building a stool), each annotated with an ordered, human-written list of steps.
ProceL [elhamifar2019unsupervised] contains 47.3 hours of annotated instructional video from 720 clips across 12 tasks (6̃0 per task). Five tasks overlap with Inria [alayrac2016unsupervised], while seven are new; videos are sourced from YouTube and trimmed for relevance.
EgoProceL [bansal2022my] comprises 62 hours of egocentric video across 16 domains, emphasizing key steps for task completion rather than every action. Key step orders may vary across videos.
@@ 4.6 Evaluation
Evaluation in procedural learning typically employs metrics like F1 score, Precision, and Recall [bansal2024united][alayrac2016unsupervised], especially in the context of error detection. Other approaches may utilize Intersection over Union (IoU) [chowdhury2024opel], the frame-level metric MoF [kukleva2019unsupervised], as well as the Hungarian algorithm, MidH, and the Jaccard index [zhou2018towards].
@@
MidH measures the percentage of predicted action segments whose midpoint falls within a ground truth action segment. Some methods employ the Jaccard score, which quantifies evaluation by computing the maximum intersection between predicted proposals and ground truth segments for each video, averaging these per video, and then averaging across all videos to obtain the final metric.
These metrics often yield high scores for models that cluster frames into a single group, since the majority key-step typically corresponds to background frames in ground truth. In untrimmed procedural videos, this background predominance can artificially inflate evaluation results. Shen et al. [shen2021learning] examined this issue using the MoF metric, which, as noted by [kukleva2019unsupervised], is unsuitable for imbalanced datasets. Bansal et al. [bansal2022my] propose computing frame-level scores per key-step and averaging over all key-steps, penalizing uneven performance, such as assigning all frames to a single key-step. This protocol results in lower scores for all methods.
@@ 5.1 Definiton
Table 5: Performance comparison of representative methods of egocentric video question answering, where metric values are averaged across different tasks within the same dataset, the bold text represents the best performance on this metric.
Method Year Accrarcy Comp. Coher. QA Score Benchmark
Lifelong (Claude-3-Haiku) [wang2023lifelongmemory] arXiv 2023 55.2 - - - EgoSchema
Lifelong (GPT-4) [wang2023lifelongmemory] arXiv 2023 62.1 - - -
Lifelong (GPT-4o) [wang2023lifelongmemory] arXiv 2023 64.6 - - -
Lifelong (Llama3-8B) [wang2023lifelongmemory] arXiv 2023 60.4 - - -
Lifelong (GPT-3.5 ) [wang2023lifelongmemory] arXiv 2023 64.0 - - -
HCQA [zhang2024hcqa] arXiv 2024 75.0 - - -
MM-Ego [ye2024mm] arXiv 2024 69.0 - - -
HCQA 1.5 [zhang2025hcqa] arXiv 2025 77.3 - - -
EgoGPT (EgoIT) [yang2025egolife] CVPR 2025 73.2 - - -
EgoGPT (EgoIT+D1) [yang2025egolife] CVPR 2025 75.4 - - -
MM-Ego [ye2024mm] arXiv 2024 61.3 - - - EgoMemoria
ALANAVLM [suglia2024alanavlm] arXiv 2024 46.7 - - - OpenEQA
GeLM [chen2025grounded] AAAI 2025 - - - 4.8 MULTIHOP-EGOQA
EgoGPT (EgoIT) [yang2025egolife] CVPR 2025 32.4 - - - EgoPlanbench
EgoGPT (EgoIT+D1) [yang2025egolife] CVPR 2025 33.4 - - -
EgoGPT (EgoIT) [yang2025egolife] CVPR 2025 61.7 - - - EgoThink
EgoGPT (EgoIT+D1) [yang2025egolife] CVPR 2025 61.4 - - -
RAVEN [biswas2025raven] arXiv 2025 54.0 - - -
EgoGPT (EgoIT) [yang2025egolife] CVPR 2025 33.1 - - - EgolifeQA
EgoGPT (EgoIT+D1) [yang2025egolife] CVPR 2025 36.0 - - -
RAVEN [biswas2025raven] arXiv 2025 61.0 71.0 69.0 - AVS-QA
Figure 7: (a) shows the different methods that using LLMs as the main inference engine to obtaining answers: Lifelongmemory [wang2023lifelongmemory], HCQA [zhang2024hcqa], HCQA 1.5 [zhang2025hcqa], FPG [taluzzi2025pixels], Egolife [yang2025egolife], and (b) shows the different methods that training specialized videoQA model to obtaining answers: GroundVQA [di2024grounded], Vinci [huang2024vinci], ALANAVLM [suglia2024alanavlm], GeLM [chen2025grounded], RAVEN [biswas2025raven].
@@ 5.2 Leveraging LLMs As Reasoning Engines
Wang et al. [wang2023lifelongmemory] introduce a question-answering framework for long-term egocentric videos utilizing Large Language Models (LLMs). The system first employs a pre-trained multimodal large model to segment videos and generate interval-based textual descriptions. These descriptions are then compressed via text summarization to remove redundancy and irrelevance. The condensed descriptions, together with user queries, are fed into an LLM to produce answers, candidate time intervals, and confidence scores. Post-processing with a pre-trained Natural Language Query (NLQ) model [ramakrishnan2023naq] further refines the temporal granularity, and iterative reasoning is used to select the answer with the highest confidence.
Zhang et al. [zhang2024hcqa] propose a three-stage pipeline (HCQA) for multiple-choice question answering on the EgoSchema egocentric video benchmark. First, the 180-second video is divided into 45 four-second clips, with the LaViLa model generating five diverse captions per clip. Second, context-driven summarization integrates these captions into a coherent global summary using GPT-4o, leveraging in-context learning for temporal comprehension. Finally, inference is performed via Chain-of-Thought reasoning to produce step-by-step answers.
Subsequently, the team introduced HCQA 1.5 [zhang2025hcqa], which enhances the original HCQA by replacing single-model inference with a dual-stage decision protocol. In the Multi-source Aggregation stage, multiple leading LLMs (e.g., Gemini-1.5-Pro, GPT-4.1, Qwen2.5) independently generate answers and confidence scores. Responses with high confidence ($\geq$4) are output directly. For lower-confidence cases, a Fine-grained Reasoning stage applies two distinct reasoning paths using DeepSeek-R1. The output with the higher confidence score is selected as the final answer, improving robustness under ambiguous visual conditions.
Taluzzi et al. [taluzzi2025pixels] introduce two graph-structured video question-answering approaches (FPG). SceneNet extracts structured scene graphs using an MLLM (Gemini 2.0 Flash) to represent objects, attributes, spatial and temporal relations, with specialized nodes for Agent, Environment, and Dynamic Objects. Graphs are generated segment-wise and temporally unified. KnowledgeNet constructs object-centric commonsense knowledge graphs by leveraging external resources (e.g., ConceptNet); Sentence Transformers perform domain filtering to retain relevant pathways, yielding structured textual outputs for enhanced reasoning.
Yang et al. [yang2025egolife] introduce EgoButler, an egocentric AI assistant comprising two subsystems: EgoGPT, a multimodal model (fine-tuned from LLaVA-OneVision) for visual and audio understanding with personalized recognition, and EgoRAG, a Retrieval-Augmented Generation module enabling long-term, cross-day information retrieval. Final responses are generated by an LLM (e.g., GPT-4o) using contextually relevant memory clips.
@@ 5.3 Training Specialized VideoQA Model
Di et al. [di2024grounded] present an end-to-end framework unifying query localization and answering to reduce error propagation. The model encodes questions and candidate answers with Flan-T5 and extracts video features via EgoVLP [lin2022egocentric] or InternVideo [chen2022internvideo], projecting them into a shared language space. Multimodal fusion is achieved with a Transformer, while a dual-head decoder predicts temporal segments and answers, supporting both open-ended and multiple-choice QA.
Huang et al. [huang2024vinci] introduce Vinci, a real-time embodied assistant based on the EgoVideo-VL model, which combines EgoVideo [pei2024egovideo] and InternLM-7B [cai2024internlm2], instruction-tuned on Ego4D, EgoExoLearn, and Ego4D-Goalstep. Vinci processes continuous egocentric video, supports real-time and historical QA, and provides task planning. It features a memory module for temporal context, a SEINE-based video generation module for visual demonstrations, and a retrieval system (using FAISS) for relevant instructional content from HowTo100M. Multimodal reasoning and answer synthesis are conducted via InternLM-7B.
Suglia et al. [suglia2024alanavlm] introduce a multimodal foundation model for egocentric video understanding, trained on the newly constructed EVUD dataset, comprising annotated and synthetic QA pairs from Ego4D NLQ [mo2022simple] and converted VSR data. Using Chat-UniVi [jin2024chat], fine-tuned with LoRA for parameter efficiency and incorporating a rehearsal mechanism to mitigate catastrophic forgetting, the model is explicitly optimized for egocentric video QA. This represents the first VLM training pipeline tailored to egocentric data, yielding robust video comprehension and QA abilities.
Chen et al. [chen2025grounded] propose GeLM, an end-to-end multimodal large language model enabling multi-hop temporal reasoning in egocentric video QA. GeLM employs InternVideo [wang2022internvideo] for video feature extraction and Vicuna-7B [chiang2023vicuna] as its language backbone. Special grounding tokens (<T><T>, </T></T>) are integrated to mark information requiring temporal localization. Its evidence localization module incorporates a saliency branch for global evidence detection and a similarity branch to align grounding tokens with relevant frames, thereby supporting precise multi-segment reasoning.
Biswas et al. [biswas2025raven] present RAVEN, a unified multimodal QA framework centered on the QuART module, which employs a learnable relevance projection head $\mathbf{W}^{R}$ to assign query-conditioned weights to each token. Tokens are fused via weighted summation to generate a context vector $C$, selectively emphasizing relevant information while suppressing noise. This design enhances robustness to misalignment and underpins a three-stage training regimen.
Ye et al. [ye2024mm] propose a two-stage progressive QA framework: “Global Glimpse” first compresses visual embeddings to obtain global context and rapidly identifies key frames; “Fallback” then uses high-resolution embeddings of those frames to generate precise answers. SigLIP-so400m extracts frame features, mapped via MLP into the Qwen2-7B LLM space. The Memory Pointer Prompting mechanism facilitates relevance-based key frame localization followed by detailed answer generation, balancing efficiency and accuracy. Training addresses both frame localization and answer generation.
@@
EVUD [suglia2024alanavlm] arXiv 2024 21.3h - 2.4W 9 10.5W Mixed Ego
@@ 5.4 Datasets
EQA [das2018embodied] is a comprehensive benchmark for embodied agent training and evaluation, built on the House3D platform with over 45,000 SUNCG-based [song2017semantic] 3D indoor scenes. It defines queries involving 12 room types and 50 object categories.
AssistQ [wong2022assistq] comprises 100 egocentric instructional videos (mean duration: 115 seconds) covering 25 household appliances, with 531 multi-step QA pairs. Questions are classified into general and specific functional categories, addressing appliance operation and settings.
EgoSchema [mangalam2023egoschema] comprises 5,063 video-question pairs drawn from 250 hours of Ego4D egocentric footage, with 3-minute clips annotated by $\geq$30 timestamps. It includes diverse question types, such as action reasoning and scene understanding. EgoSchema introduces the Temporal Certificate metric, which represents the minimal cumulative duration of segments required to verify annotation validity.
VIEW-QA [song2024video] comprises 1,030 videos (1̃0 hours) and 4,120 QA pairs, with questions divided into five categories reflecting the practical needs of visually impaired users. Its dynamic, multi-task video QA design better mirrors real-world use cases.
EVUD [suglia2024alanavlm] supports both video captioning and QA tasks, comprising 13,849 Ego4D NLQ clips with 1,137 human-annotated QA pairs, 96,523 Gemini Pro 1.5-generated QA pairs across 7 categories, 7,680 polar QA pairs for spatial reasoning, 7,000 EgoClip videos with captions, and 3,475 short videos with captions generated via Habitat for HM3D scenes in OpenEQA [majumdar2024openeqa].
HOI-QA [bansal2024hoi] comprises 3.9 million QA pairs: 1.8M from EPIC-Kitchens (with narratives, object bounding boxes, and hand-object contact) and 2.1M from Ego4D’s FHO benchmark (narratives, hand/object bounding boxes). The data is split into 3.2M training samples and 0.75M test samples.
EgoTempo [plizzari2025omnia] features 500 QA pairs spanning 10 temporal reasoning tasks using egocentric videos (avg. 45 seconds) from 40 scenarios, designed to evaluate MLLM temporal integration abilities.
EgoBlind [xiao2025egoblind] is the first QA dataset constructed from real-world egocentric videos recorded by blind individuals. Unlike prior datasets such as VizWiz (static images) [gurari2018vizwiz] and VIEW-QA (simulated blindness) [song2024video], EgoBlind uniquely addresses the dynamic informational needs of blind users, which are not captured by general video QA resources.
EgoTextVQA [zhou2025egotextvqa] targets scene text understanding in egocentric contexts, comprising 1,507 curated indoor and outdoor videos and 7,064 QA pairs. Videos were filtered using OCR [li2022pp] from RoadTextVQA [tom2023reading], EgoSchema, and Ego4D, with manual quality control to ensure the presence of relevant text.

### v2

@@ 4.5 Datasets
For procedural error detection, various egocentric datasets have been introduced, covering scenarios such as cooking [damen2020epic][peddi2024captaincook4d], assembly [sener2022assembly101][liu2024ikea], outdoor tent pitching [jang2019epic], and specialized domains like chemical experiments [haneji2025egooops]. Several additional datasets are also prevalent in procedural learning research, as shown in Table 4.
Breakfast [kuehne2014language] includes 77 hours of video for 10 cooking activities by 52 participants in 18 kitchens. Each video averages 5.2 unique actions and 6.9 action segments, with 10% of actions repeated. Inria [alayrac2016unsupervised] is an instructional dataset of 150 videos (30 per task) spanning five tasks, with 800,000 frames and comprehensive annotations. YouCook2 [zhou2018towards] contains 2,000 untrimmed videos featuring 89 global cooking recipes, averaging 22 videos per recipe and encompassing diverse cooking styles. COIN [tang2019coin] contains 11,827 YouTube videos across 180 tasks in 12 daily domains, totaling 476 hours and 46,354 annotated segments, with an average video length of 2.36 minutes. HowTo100M [miech2019howto100m] includes 1.22 million instructional videos, covering 23,000+ visual tasks and 136 million clips, each with narrations and auto-downloaded subtitles from YouTube. CrossTask [zhukov2019cross] consists of instructional videos for 83 tasks (e.g., making a bread, preparing a latte, building a stool), each annotated with an ordered, human-written list of steps. ProceL [elhamifar2019unsupervised] contains 47.3 hours of annotated instructional video from 720 clips across 12 tasks (6̃0 per task). Five tasks overlap with Inria [alayrac2016unsupervised], while seven are new; videos are sourced from YouTube and trimmed for relevance. EgoProceL [bansal2022my] comprises 62 hours of egocentric video across 16 domains, emphasizing key steps for task completion rather than every action. Key step orders may vary across videos.
@@ 4.6 Evaluation
Evaluation in procedural learning typically employs metrics such as F1 score, Precision, and Recall [bansal2024united][alayrac2016unsupervised], especially for error detection. Other approaches may utilize Intersection over Union (IoU) [chowdhury2024opel], the frame-level metric MoF [kukleva2019unsupervised], the Hungarian algorithm, MidH, and the Jaccard index [zhou2018towards].
@@
MidH measures the percentage of predicted action segments whose midpoints fall within a ground-truth action segment. Some methods employ the Jaccard score, which quantifies evaluation by computing the maximum intersection between predicted proposals and ground truth segments for each video, averaging these per video, and then averaging across all videos to obtain the final metric.
These metrics often yield high scores for models that cluster frames into a single group, since the majority key-step typically corresponds to background frames in ground truth. In untrimmed procedural videos, this background predominance can artificially inflate evaluation results. Shen et al. [shen2021learning] examined this issue using the MoF metric, which, as noted by [kukleva2019unsupervised], is unsuitable for imbalanced datasets. Bansal et al. [bansal2022my] propose computing frame-level scores per key-step and averaging over all key-steps, penalizing uneven performance, such as assigning all frames to a single key-step. This protocol yields lower scores across all methods.
@@ 5.1 Definiton
Table 5: Performance comparison of representative methods of egocentric video question answering, where metric values are averaged across different tasks within the same dataset. The bold text represents the best performance on this metric.
Method Year Accuracy QA Score (0-10) Benchmark
Lifelong (Claude-3-Haiku) [wang2023lifelongmemory] arXiv 2023 55.2 - EgoSchema
Lifelong (GPT-4) [wang2023lifelongmemory] arXiv 2023 62.1 -
Lifelong (GPT-4o) [wang2023lifelongmemory] arXiv 2023 64.6 -
Lifelong (Llama3-8B) [wang2023lifelongmemory] arXiv 2023 60.4 -
Lifelong (GPT-3.5 ) [wang2023lifelongmemory] arXiv 2023 64.0 -
HCQA [zhang2024hcqa] arXiv 2024 75.0 -
MM-Ego [yemmego] ICLR 2025 69.0 -
HCQA 1.5 [zhang2025hcqa] arXiv 2025 77.3 -
EgoGPT (EgoIT) [yang2025egolife] CVPR 2025 73.2 -
EgoGPT (EgoIT+D1) [yang2025egolife] CVPR 2025 75.4 -
LLaVA-OneVision [lillava] TMLR 2025 47.3 - EgoMemoria
MM-Ego [yemmego] ICLR 2025 61.3 -
Chat-UniVi [jin2024chat] CVPR 2024 42.3 - OpenEQA
ALANAVLM [suglia2024alanavlm] EMNLP 2024 46.7 -
GPT-4o [hurst2024gpt] arXiv 2024 - 5.4 MULTIHOP-EGOQA
GeLM [chen2025grounded] AAAI 2025 - 4.8
EgoGPT (EgoIT) [yang2025egolife] CVPR 2025 32.4 - EgoPlanbench
EgoGPT (EgoIT+D1) [yang2025egolife] CVPR 2025 33.4 -
EgoGPT (EgoIT) [yang2025egolife] CVPR 2025 61.7 - EgoThink
EgoGPT (EgoIT+D1) [yang2025egolife] CVPR 2025 61.4 -
RAVEN [biswas2025raven] arXiv 2025 54.0 -
EgoGPT (EgoIT) [yang2025egolife] CVPR 2025 33.1 - EgolifeQA
EgoGPT (EgoIT+D1) [yang2025egolife] CVPR 2025 36.0 -
Figure 6: (a) shows the different methods that using LLMs as the main inference engine to obtaining answers: Lifelongmemory [wang2023lifelongmemory], HCQA [zhang2024hcqa], HCQA 1.5 [zhang2025hcqa], FPG [taluzzi2025pixels], Egolife [yang2025egolife], and (b) shows the different methods that training specialized videoQA model to obtaining answers: GroundVQA [di2024grounded], Vinci [huang2025vinci], ALANAVLM [suglia2024alanavlm], GeLM [chen2025grounded], RAVEN [biswas2025raven].
@@ 5.2 Leveraging LLMs As Reasoning Engines
More suitable for solving egocentric long‑video question answering tasks, Wang et al. (Lifelongmemory) [wang2023lifelongmemory] and Zhang et al. [zhang2024hcqa, zhang2025hcqa] (HCQA/HCQA 1.5) are building on the powerful reasoning and descriptive capabilities of MLLMs, they generate textual descriptions of videos and leverage techniques such as context summarization and temporal localization, further incorporating in-context learning and reflection mechanisms, thereby achieving very high accuracy. Besides, Yang et al. [yang2025egolife] (Egolife) even works on week-long videos. Unlike these methods, Taluzzi et al. [taluzzi2025pixels] (FPG) construct graphs such as SceneNet and KnowledgeNet to obtain powerful reasoning and memory capabilities. It performs well on classification tasks and supports cross-modal reasoning, making it more suitable for relational analysis tasks.
@@ 5.3 Training Specialized VideoQA Model
Chen et al. (GeLM) [chen2025grounded] propose an end-to-end multimodal large language model which employs grounded tokens (<T><T>, </T></T>) together with a dual-branch (saliency + similarity) evidence localization module. Its limitations lie in the reliance on manually curated multi-hop annotations and the insufficient handling of cross-modal ambiguities. With lower data annotation cost, Di et al. (GroundVQA) [di2024grounded] achieves multimodal fusion with a Transformer, while a dual-head decoder predicts temporal segments and answers, supporting both open-ended and multiple-choice QA. Also, with low data requirements, Suglia et al. (ALANAVLM) [suglia2024alanavlm] apply LoRA-based efficient fine-tuning to Chat-UniVi [jin2024chat], focusing on embedded scenario question answering. The approach yields a lightweight model (7B parameters) whose spatial reasoning performance surpasses that of GPT-4V. Fuse video, audio, and sensor modalities, Biswas et al. (RAVEN) [biswas2025raven] dynamically filters irrelevant tokens via the QuART module. Combined with a three-stage training scheme, it exhibits strong cross-modal robustness. However, frame sampling in long videos often misses critical information. To better address long video processing, Ye et al. (MM-EGO) [yemmego] rely on a memory-pointer prompting mechanism (global overview + key-frame fallback), achieving high efficiency in long-video processing and strong robustness against language bias. Huang et al. (Vinci) [huang2025vinci] introduce a real-time intelligent assistant built on the EgoVideo-VL model, integrating a memory module and video generation and retrieval modules. It supports audio-based interaction and step-by-step visual demonstrations, offers strong real-time performance, and enables effective use of historical context and visually guided generation.
@@
EVUD [suglia2024alanavlm] EMNLP 2024 21.3h - 2.4W 9 10.5W Mixed Ego
@@ 5.4 Datasets
EQA [das2018embodied] is a comprehensive benchmark for embodied agent training and evaluation, built on the House3D platform with over 45,000 SUNCG-based [song2017semantic] 3D indoor scenes. It defines queries involving 12 room types and 50 object categories. AssistQ [wong2022assistq] comprises 100 egocentric instructional videos (mean duration: 115 seconds) covering 25 household appliances, with 531 multi-step QA pairs. Questions are classified into general and specific functional categories, addressing appliance operation and settings. EgoSchema [mangalam2023egoschema] comprises 5,063 video-question pairs drawn from 250 hours of Ego4D egocentric footage, with 3-minute clips annotated by $\geq$30 timestamps. It includes diverse question types, such as action reasoning and scene understanding. EgoSchema introduces the Temporal Certificate metric, which represents the minimal cumulative duration of segments required to verify annotation validity. VIEW-QA [song2024video] comprises 1,030 videos (1̃0 hours) and 4,120 QA pairs, with questions divided into five categories reflecting the practical needs of visually impaired users. Its dynamic, multi-task video QA design better mirrors real-world use cases. EVUD [suglia2024alanavlm] supports both video captioning and QA tasks, comprising 13,849 Ego4D NLQ clips with 1,137 human-annotated QA pairs, 96,523 Gemini Pro 1.5-generated QA pairs across 7 categories, 7,680 polar QA pairs for spatial reasoning, 7,000 EgoClip videos with captions, and 3,475 short videos with captions generated via Habitat for HM3D scenes in OpenEQA [majumdar2024openeqa]. HOI-QA [bansal2024hoi] comprises 3.9 million QA pairs: 1.8M from EPIC-Kitchens (with narratives, object bounding boxes, and hand-object contact) and 2.1M from Ego4D’s FHO benchmark (narratives, hand/object bounding boxes). The data is split into 3.2M training samples and 0.75M test samples. EgoTempo [plizzari2025omnia] features 500 QA pairs spanning 10 temporal reasoning tasks using egocentric videos (avg. 45 seconds) from 40 scenarios, designed to evaluate MLLM temporal integration abilities. EgoBlind [xiao2025egoblind] is the first QA dataset constructed from real-world egocentric videos recorded by blind individuals. Unlike prior datasets such as VizWiz (static images) [gurari2018vizwiz] and VIEW-QA (simulated blindness) [song2024video], EgoBlind uniquely addresses the dynamic informational needs of blind users, which are not captured by general video QA resources. EgoTextVQA [zhou2025egotextvqa] targets scene text understanding in egocentric contexts, comprising 1,507 curated indoor and outdoor videos and 7,064 QA pairs. Videos were filtered using OCR [li2022pp] from RoadTextVQA [tom2023reading], EgoSchema, and Ego4D, with manual quality control to ensure the presence of relevant text.